### v1

# Binding Affinity Prediction: From Conventional to Machine Learning-Based Approaches

Abstract
Protein-ligand binding is the process by which a small molecule (drug or inhibitor) attaches to a target protein. The binding affinity, which refers to the strength of this interaction, is central to many important problems in bioinformatics such as drug design. An extensive amount of work has been devoted to predicting binding affinity over the past decades due to its significance. In this paper, we review all significant recent works, focusing on the methods, features, and benchmark datasets. We have observed a rising trend in the use of traditional machine learning and deep learning models for predicting binding affinity, accompanied by an increasing amount of data on proteins and small drug-like molecules. While prediction results are constantly improving, we also identify several open questions and potential directions that remain unexplored in the field. This paper could serve as an excellent starting point for machine learning researchers who wish to engage in the study of binding affinity, or for anyone with general interests in machine learning, drug discovery, and bioinformatics.
1 Introduction
Protein-ligand binding (Clyde et al., 2023) refers to the process as shown in Fig. 1 by which ligands—usually small molecules, ions, or proteins—generate signals by binding to the active sites of target proteins through intermolecular forces. This binding typically changes the conformation of target proteins, which then results in the realization, modulation, or alteration of protein functions. Therefore, protein-ligand binding plays a central role in most, if not all, important life processes. For example, oxygen molecules are bound and carried through the human body by proteins like hemoglobin, and then utilized for energy production, while nonsteroidal anti-inflammatory drugs (NSAIDs) like ibuprofen work by inhibiting the functionality of the cyclooxygenase (COX) enzyme that thus reducing the release of pain-causing substances in the body.
The concept and importance of binding affinity prediction were first addressed in Böhm (1994): given the 3D structures of a target protein and a potential ligand, the objective is to predict the binding constant of such a complex, along with the most probable binding pose candidates. The prediction of the binding site (the set of protein residues that have at least one non-hydrogen atom within 4.0 Å of a ligand’s non-hydrogen atom (Khazanov and Carlson, 2013)) and affinity (binding constants such as inhibition or dissociation constants, or the concentration at 50% inhibition) are usually divided into two separate but related stages (Ballester and Mitchell, 2010a).
Figure 1: Binding Example in Surface View with Protein PDB: 10GS
One notable motivation for constructing a good binding affinity predictor (or scoring function, as called in some earlier work) is the essential role that it plays in drug discovery (Liu et al., 2023, 2024a) and virtual screening (Meng et al., 2011; Pinzi and Rastelli, 2019; Sadybekov and Katritch, 2023). Traditional drug discovery essentially involves a process of trial and error. However, with a functional binding affinity predictor, we can significantly reduce the number of experimental trials by focusing only on drugs estimated to be more effective, thus potentially saving large amounts of both time and money. This motivation is becoming increasingly relevant with the rising costs of drug discovery (Scannell et al., 2012; Hay et al., 2014; Tiwari et al., 2023) and the declining rate of drug approvals over recent decades (Dickson and Gagnon, 2004; Paul et al., 2010). Beyond its economic significance in drug development, binding affinity prediction also provides tools, methods, and insights into many other areas of research in biochemistry, pharmaceutical research, and scientific computation.
Despite the increasing motivation and extensive amount of work, binding affinity prediction remains challenging. First and foremost, our understanding of chemistry is not nearly as sufficient as it needs to be, which often leads to sub-optimal human-engineered features for binding affinity prediction, as shown in Ballester and Mitchell (2010a) and Ballester et al. (2014). Furthermore, synthetic datasets remain undesirable for learning, because computationally generated complexes are either inaccurate or too costly to generate in terms of time and computational resources.
Currently, the datasets used, which are usually obtained experimentally for the aforementioned reason, are each limited in various ways: (1) the total number of data points is increasing but still far from sufficient for large-scale data mining due to the vast time and money required to study each complex; (2) the accuracy of experimental data may be limited by the measurement methods used (Su et al., 2018; Pantsar and Poso, 2018); and (3) samples are often biased toward complexes that have the correct poses and good binding constants, meaning that protein-ligand pairs that fail to bind or have relatively low binding constants are not readily obtainable and therefore are unavailable for learning. All these limitations make the search for a good protein-ligand binding affinity predictor extremely challenging.
To make things more complicated, there are different aspects/sub-problems associated with the prediction of binding constants. Li et al. (2014) and Liu et al. (2017) note four different types of predictions related to binding affinity: (1) scoring, which predicts the binding constant; (2) ranking, which ranks different potential ligands of a target; (3) docking, which predicts the best binding pose from multiple decoys; and (4) screening, which predicts the best ligand from multiple decoys. These sub-problems are closely related to each other, to the extent that solving one without addressing the others adds to the existing challenges. Thus, despite all the work and effort put into protein-ligand binding affinity prediction, the results remain sub-optimal. It is often the case that a predictor excels in scoring but fails miserably in docking or screening, rendering the whole system almost useless for virtual screening, one of the ultimate goals of binding affinity prediction.
Figure 2: Binding affinity pipeline (in cartoon view): Computational binding affinity measurements are typically performed using molecular docking simulations. The simulations search for an optimal binding pose for each compound and produces a score for each compound. These scores are then used to rank compounds with respect to each other.
Nevertheless, all previous works can be roughly divided into three distinct categories based on the methods used for binding affinity prediction: conventional, traditional machine-learning-based, and deep-learning-based. Conventional methods, as shown in Fig. 2, are typically based purely on knowledge and assumptions about chemistry and quantum mechanics, often formulated as a set of rules or equations that combine different factors into a single binding affinity result. These methods are usually rigid and tend to work well only in specific scenarios, such as with a single protein family (Bender et al., 2021).
Since 2005, many traditional machine learning methods (such as random forest and support vector machines) have been applied to human-engineered features extracted from complex structures, achieving some improvement over conventional approaches. Machine learning methods have proven to be less rigid and often more accurate, especially in terms of binding affinity scoring and ranking (Ballester and Mitchell, 2010b; Zilian and Sotriffer, 2013; Li et al., 2011) A decade later, as the number of protein-ligand samples in standard benchmarks like PDBbind (Wang et al., 2004a), MOAD (Hu et al., 2005), and PDB (Burley et al., 2017) increased, deep learning began to dominate. This approach typically relies less on human-engineered features, if at all, and its learning potential could greatly increase with the continued accumulation of data.
This paper covers important work on protein-ligand binding affinity prediction, from the early 2000s to the present, along with commonly used datasets and evaluation metrics. The remainder of the paper is organized as follows: Section 2 introduces important datasets and benchmarks, including data specifications and various datasets. Section 3 discusses methodology, in two categories: conventional and machine learning-based methods. In Section 4, we introduce various evaluation approaches, including scoring power, docking power, and ranking power. Finally, we conclude with a discussion of open questions and possible directions that remain unexplored in the field.
2 Datasets and Benchmarks
There are many datasets and benchmarks used for the study of protein-ligand binding affinity prediction, each focusing on different aspects of the problem. Here, we first discuss in Section 2.1 the specifications for binding affinity data, such as resolution, measurement, binding constant, and concentration. These factors are essential when selecting the right dataset for a specific goal. In Section 2.2, we review some of the most commonly used datasets and benchmarks directly applied in the learning and prediction of protein-ligand binding affinity. These datasets usually have binding constants such as $K_{i}$, $K_{d}$, or $IC_{50}$ associated with most, if not all, protein-ligand complexes. We also note other datasets of protein-ligand binding affinity not directly used for model training and evaluation. Even though these datasets are not directly applicable for training and evaluation, they can be useful for pre-training and may provide additional information, methods, and insights. All discussed datasets are listed in Table 1.
Name # Proteins # Ligands # Complexes # Affinities 3D Primary Sources
PDBbind 9,198 13,427 19,588 19,588 Yes PDB
CASF 57 285 285 285 Yes PDB
Binding MOAD 9,117 16,044 32,747 12,101 Yes PDB
BindingDB 7,317 751,447 1,692,135 1,692,135 Partial Publications, PubChem, ChEMBL
DUD-E 102 22,886 22,886 22,886 Yes PDB, ChEMBL, ZINC
BioLiP 97,966 460,356 460,364 23,492 Yes Binding MOAD, PDBbind, BindingDB, Pubications
PDSP Ki 958 12,228 67,689 67,689 No Publications
KIBA 467 52,498 246,088 246,088 No ChEMBL
Table 1: Commonly Used Protein-Ligand Binding Affinity Datasets
2.1 Data Specifications
2.1.1 Structure
The structure of a protein-ligand complex refers to its conformation, or in other words, the spatial arrangement of its atoms. This is an important specification or feature of binding affinity data, as it opens up numerous possibilities for feature engineering and learning methods. For instance, with given structures, we can parameterize the energy terms and atom-pair distances and use them for conventional scoring or learning with methods such as random forests (Ballester and Mitchell, 2010a), support vector machines (Kinnings et al., 2011), and neural networks (Durrant and McCammon, 2010, 2011). On the other hand, measuring structures is rather expensive (more than \$1,000 USD for solving a structure by crystal diffraction) compared to sequence identity, along with the additional costs of data storage and manipulation. Still, the advantages of having a well-measured structure significantly outweigh the disadvantages, and we are observing a rising trend of more protein-ligand complexes with structures being added to PDB (Protein Data Bank) and other databases.
2.1.2 Resolution
Resolution, in the context of protein-ligand complex structures, refers to the distance corresponding to the smallest observable feature in the measured structure (Guterres and Im, 2020). In other words, two objects within this distance will be perceived as one and thus rendered indistinguishable. Generally, structures with a resolution finer than 1–1.2 Å are considered high-resolution, while those with a resolution lower than 3.0 Å only outline the basic contours of the protein chains. Resolution is an essential measurement for the quality of data and is used as a selection criterion for many binding affinity databases, such as CASF and Binding MOAD.
2.1.3 Structure Determination Techniques
The three most commonly used techniques for determining the structure of protein complexes are X-ray crystallography (Maveyraud and Mourey, 2020; Jackson et al., 2023), NMR (nuclear magnetic resonance) (Hore, 2015; Galvan et al., 2023), and Cryo-EM (cryo-electron microscopy) (Adrian et al., 1984; Nogales and Mahamid, 2024).
Roughly speaking, X-ray crystallography determines the position and arrangement of atoms in a single crystal of the target protein by examining the diffraction intensity obtained with X-rays. Given high-quality protein crystal, X-ray crystallography is capable of generating atom-level static structures, regardless of molecular weight of the samples.
NMR is a fundamentally different technique for protein structure determination. The analysis is performed on a solution of the target protein with high purity and high concentration to obtain characteristic NMR signals, which are interpreted by computer-aided methods to determine the 3D structures. The most notable feature of NMR is that it allows us to obtain the dynamic structure of the target protein in its natural state in solution (without crystallization). However, the measurement is complicated, requires computational interpretation (making NMR an indirect method of structure determination), and is not applicable to large molecules or samples without pure and highly-concentrated solutions.
Cryo-EM uses a mechanism called electron scattering, where electron beams pass through an instantly cooled protein solution, scatter into a lens, and are then converted into a series of 2D images on a detector. These images are processed by reconstruction software to obtain the 3D structure. Cryo-EM is becoming increasingly popular because it works well with larger proteins, has an easier sample preparation process, and yields 3D structures that are much closer to their native state than those obtained from X-ray crystallography. However, the resolution of the structures can be relatively low due to the high level of noise and unknown orientations.
2.1.4 Binding Affinity Measurement
Roughly speaking, binding affinity measures how “tightly” a ligand binds to a target protein, which is determined with radioligand-binding experiments (Haylett, 2003). A radioligand-binding experiment can produce two kinds of binding affinity data: (1) binding constants such as dissociation constant ($K_{d}$) and inhibition constant ($K_{i}$), which are usually interchangeable in the context of protein-ligand binding; and (2) concentration terms like the half maximal inhibitory concentration (${IC}_{50}$) or half maximal effective concentration (${EC}_{50}$). In binding affinity databases, each protein-ligand complex is typically associated with one or more binding data of $K_{d}$, $K_{i}$, and ${IC}_{50}$, with rare exception. In certain databases and studies, like the “refined set” in PDBbind, complexes with concentration terms only will be filtered out for better data quality. This is due to the fact that ${IC}_{50}$ or ${EC}_{50}$ values depend on the radioligand concentration and can vary between different experiments, unlike constant terms. If necessary or useful, the inhibitory constant $K_{i}$ value can be estimated from $IC_{50}$ via the Cheng-Prusoff equations:
$K_{i}=\frac{\text{IC}_{50}}{\frac{[A]}{\text{EC}_{50}}+1}\;\text{,}$
where $[A]$ is a fixed concentration of the ligand. However, the estimation is prone to inaccuracy with high or low concentration values (Lazareno and Birdsall, 1993).
2.2 Binding Affinity Datasets
PDBbind. Starting in 2004, this dataset has been updated annually (Wang et al., 2005; Liu et al., 2015, 2017). It collects not only protein-ligand complexes but also protein-protein, protein-nucleic acid, and nucleic acid-ligand complexes from the Protein Data Bank (PDB) (Berman et al., 2000), without any restrictions on resolution, binding data, or structural measurement techniques. An important feature of PDBbind is the refined set, selected based on (1) the quality of the complex structure, (2) the quality of binding data, and (3) the nature of the complex (molecular weight, atom types, surface area, etc.). See the supporting material in (Liu et al., 2017) for more details on the selection criteria. However, according to the authors, the refined set, despite its much higher quality compared to the general set, should not be regarded as a high-quality dataset but rather as a collection of complex samples that lack any obvious problems. With the increasing number of data points, subsets with quality control, and easy accessibility, PDBbind is becoming one of the most used datasets in the research of protein-ligand binding affinity, if not the most used.
CASF (Comparative Assessment of Scoring Functions). CASF (Cheng et al., 2009; Li et al., 2018; Su et al., 2018) was proposed along with the development of PDBbind. Originally named the core set in PDBbind, CASF was selected from the refined set and serves as the test set/benchmark for all the protein-ligand binding predictor. To ensure that the samples in the test set are diverse but not redundant, all complexes in the refined set are first clustered based on sequence similarity. Then, for each cluster, three complexes, each with high, medium, and low binding affinity, are selected for inclusion in CASF. With this type of data selection, we can not only test regression on binding affinity (scoring) but also ranking (three different complexes in the same cluster). Additionally, with the given decoys of poses and ligands in CASF, we can evaluate the docking and screening power of a predictor. Overall, CASF provides researchers with a comprehensive and easy-to-use benchmark for protein-ligand binding affinity predictors.
Binding MOAD (Mother of All Databases). Binding MOAD (Hu et al., 2005; Ahmed et al., 2015; Smith et al., 2019) is another commonly used dataset for binding studies. Proposed in 2005 and updates annually, Binding MOAD is a subset of PDB, aiming to be the largest collection of high-quality protein-ligand complexes annotated with experimentally determined binding affinity. Despite sharing the same primary data source (PDB) and a similar goal, Binding MOAD has much in common with PDB. The main differences that set these two apart for users are: (1) Binding MOAD sets the data selection criteria somewhere between PDBbind’s general set and refined set. Notably, Binding MOAD only contains valid protein-ligand complexes with crystal structures of 2.5 Åresolution or better. Additionally, protein sequences are clustered to avoid redundancy of the data, similar to the refined set in PDBbind. (2) Binding MOAD contains complexes without binding data, which is not the case in PDBbind. In fact, Binding MOAD contains only 12,098 binding data of 32,747 complexes. This means that binding MOAD, while having more complexes with structures than PDBbind, contains fewer samples with binding data.
BindingDB. Launched in 2000 and updated weekly as a web-accessible database, BindingDB (Liu et al., 2007; Gilson et al., 2016) collects protein-compound pairs with associated affinity data primarily from scientific articles and, increasingly, patents. It also accepts direct deposition of binding data by users. Newly curated data from these original sources are checked by BindingDB staff to ensure reliability. Additionally, it integrates complexes with affinity measurements from PubChem (Wang et al., 2009), ChEMBL (Bento et al., 2014), PDSP Ki (Roth et al., 2000), and CSAR (Carlson and Dunbar Jr, 2011), allowing users to access all the data through the unified interface of BindingDB (Gilson et al., 2016). There are two important differences between BindingDB and other large binding affinity datasets such as PDBbind and Binding MOAD. First, although most binding data are experimentally determined, BindingDB contains some complexes that are computationally generated with clear labels. Secondly and more importantly, only a fraction of BindingDB’s affinity data has associated structures. In fact, according to their website, there are 2,291 protein-ligand crystal structures available in BindingDB with 100% sequence identity and 5,816 crystal structures with protein sequence identity as low as 85%. The variety of data sources and the aforementioned differences between BindingDB and other databases make it the largest protein-ligand binding affinity database in terms of the number of complexes.
DUD-E (A Directory of Useful Decoys: Enhanced). To combat data bias and benchmark ligand enrichment against challenging decoys, DUD (Huang et al., 2006) (Directory of Useful Decoys) and its successor DUD-E (Mysinger et al., 2012) (A Directory of Useful Decoys: Enhanced) were proposed in 2006 and 2012, respectively. The dataset was constructed by sourcing ligands from ChEMBL, structural data from PDB, and potential decoys from ZINC. The docking poses were generated with DOCK (Ewing et al., 2001; Allen et al., 2015). Immediately after being proposed and constructed, DUD became the gold standard for the evaluation of virtual screening methods (Réau et al., 2018). However, as pointed out in Good and Oprea (2008), the original DUD dataset had a number of issues, such as target selection and analogue bias. To address these issues, the authors of DUD released DUD-E, which added net charge into consideration during decoy selection. In DUD-E, each of the 102 target proteins has 50 decoys that are similar to the ligand in terms of 1D physico-chemical properties, such as molecular weight and calculated LogP, to remove dataset bias, but differ in 2D topology to be likely non-binders. Nevertheless, DUD-E remains unsatisfactory, as shown in multiple works (Xia et al., 2015; Chen et al., 2019; Sieg et al., 2019), indicating that hidden bias in the dataset might still exist, leading to misleadingly good benchmarking results for many docking methods. A good experiment to reveal this bias is to featurize the decoy ligands and plot them using 2D or 3D t-SNE to see if the decoy and real ligands separate.
BioLiP (Yang et al., 2013a) and its successor BioLiP2 (Zhang et al., 2024) (collectively referred to as BioLiP) are semi-manually curated databases for protein-ligand interactions. Founded in 2013 and updated weekly, BioLiP is designed to focus only on biologically relevant ligands, meaning they are not simply additives for protein purification and/or crystallization. This is achieved through automatic procedures and manual verification. In addition to annotations of binding affinity, the dataset also contains relevant functional annotations of the complexes. It sources structures from PDB and binding data from multiple databases such as Binding MOAD, PDBbind, BindingDB, and literature. It is worth noting that the main focus of BioLiP is binding site prediction, similar to FireDB (Maietta et al., 2014) and LigASite (Dessailly et al., 2007), and binding affinity is one of the optional fields for each entry in the database. Only about 5% of the protein-ligand complexes have associated binding affinity.
PDSP (Psychoactive Drug Screening Program) Ki Database. The PDSP Ki Database (Roth et al., 2000), founded in 2006, serves as a regularly updated data warehouse for published binding affinity data of drugs and drug candidates for receptors, ion channels, transporters, and enzymes. The entries are accessible for query under given conditions or can be downloaded as a CSV file. For each entry, the IDs of the ligand and receptors are provided, along with additional information such as the SMILES string and the reference for the affinity data. Despite the large quantity of affinity data, there are two major drawbacks to the PDSP Ki Database. First, the affinity data ($K_{i}$ value) for each compound-receptor pair may be either a single numeric value or an inequality such as “greater than 7000.” Secondly, structure features for both protein and ligand are not included in the dataset, meaning users will have to query other data sources such as PDB.
KIBA (Kinase Inhibitor Bioactivity) Dataset was originally proposed in Tang et al. (2014). It aimed to integrate various bioactivity types, including $K_{i}$, $K_{d}$, and $IC_{50}$, into a single term named the KIBA score, and to demonstrate its usage in classifying kinase inhibitor targets and pinpointing errors in binding affinity datasets. The dataset was obtained from searching for drug-kinase interactions in ChEMBL with $IC_{50}$ and at least one binding constant terms ($K_{i}$ and $K_{d}$). For each interaction in the KIBA dataset, they have converted the bioactivity terms into KIBA score using the equations below:
${K_{i}}_{\cdot adj}$ $=\frac{{IC}_{50}}{1+L_{i}({IC}_{50}/K_{i})},~{}{K_{d}}_{\cdot adj}=\frac{{IC}_{50}}{1+L_{d}({IC}_{50}/K_{d})}$ (1)
KIBA =$=\begin{cases}{K_{i}}_{\cdot adj}&\text{if ${IC}_{50}$ and $K_{i}$ are present}\\ {K_{d}}_{\cdot adj}&\text{if ${IC}_{50}$ and $K_{d}$ are present}\\ ({K_{i}}_{\cdot adj}+{K_{d}}_{\cdot adj})/2&\text{if ${IC}_{50}$, $K_{i}$, and$K_{d}$ are present}\end{cases}$ (2)
with $L_{i}=0.3$ and $L_{d}=1.3$, which are the optimal values for adjustment. There are no features of any kind included in the original dataset. Users need to look up ChEMBL, PDB, or other sources to obtain features of drug-kinase pairs.
Other Binding Affinity Datasets. AffinDB (Block et al., 2006) was a major binding affinity dataset when it was proposed in 2006. However, the website and affinity data have not been updated for some time. As of 2019, there are 748 affinity values covering 474 PDB complexes, which is severely insufficient compared to PDBbind or BindingDB. CSAR (Community Structure-Activity Resource) (Smith et al., 2011; Dunbar et al., 2011; Damm-Ganamet et al., 2013; Dunbar et al., 2013; Smith et al., 2016; Carlson et al., 2016) was an experimental dataset of crystal structures and binding affinities for diverse protein-ligand complexes. The dataset was intended for a community-wide exercise conducted by a group at the University of Michigan from 2010 to 2014, aiming to use unpublished data from in-house projects to evaluate existing protein-ligand binding affinity prediction methods. The CSAR dataset has inspired some of the most meaningful discussions on the methods and evaluations of binding affinity, such as those in (Novikov et al., 2011; Koes et al., 2013; Carlson, 2016), and the ideas and insights from these discussions are still relevant today. Aside from those datasets that are no longer maintained, there are many datasets that are relevant to binding affinity in one aspect or another. For instance, Tang et al. (2014) cite studies of bioactivity profiling of small-molecule protein kinase inhibitor by Davis et al. (2011), Metz et al. (2011), and Anastassiadis et al. (2011), each of which provides a set of measurements of binding affinity between kinase and small compounds. Krivák and Hoksza (2018) combined multiple small protein-ligand complex datasets, such as CHEN11 (Chen et al., 2011), COACH420 (Roy et al., 2012; Yang et al., 2013b), and HOLO4K (Schmidtke et al., 2010), for training and evaluation of binding site prediction. As existing datasets and models fail to account for the dynamic features of protein-ligand interactions, the recently released PLAS-20K (2024) addresses this gap. As an extension of PLAS-5K (2022), PLAS-20K includes 97,500 independent simulations across 19,500 different protein-ligand complexes.
3 Methodology
Next, we discuss the various protein-ligand binding affinity prediction methodologies as developed from the early 2000s to the present day. We divided all the methods into two categories based on the model complexity and input data. In roughly chronological order, they are: (1) Conventional methods, which are essentially a set of equations derived from assumptions and understanding of the binding process, combining weighted physio-chemical terms in an additive manner into a single estimate for affinity. (2) Machine learning models (e.g., random forest, support vector machine, neural network, etc.) trained with human-engineered descriptors extracted from the protein-ligand complexes; and representation-learning methods, which extract features directly from raw data of protein-ligand complexes (e.g., SMILES strings, voxels, graphs) using the learning capacity of deep neural networks. Roughly speaking, as the study of binding affinity advances, the community has shifted from methods that require extensive domain knowledge and assumptions to those that can exploit the increasing number of available protein-ligand complex structures.
3.1 Conventional Methods
Ever since the formalization of the concept of protein-ligand docking, researchers have been trying to predict the binding affinity or energy based on human understanding of physics and chemistry. These methods were often called scoring functions, rather than predictors or models, because they almost always tool the forms of additive functions, combining various engineered physio-chemical terms. Traditionally, these conventional scoring functions have been roughly divided into the following three categories, in multiple works (Cheng et al., 2009; Ballester and Mitchell, 2010a; Ashtawy and Mahapatra, 2012):
3.1.1 Physics-based methods
Traditionally, these methods are often referred as force-field-based methods, because “classical force-field-based scoring functions assess the binding energy by calculating the sum of the non-bonded interactions” (Meng et al., 2011), which is often done with force field. However, Liu and Wang (2015) suggested the change of name from force-field methods to physics-based methods because some of the methods or models commonly used in this category of scoring functions, such as Poisson-Boltzmann or Generalized Born continuum solvation models, end-point approximation, and quantum mechanics computation, are not based on force fields. The new name, physics-based methods, more appropriately reflects the theoretical basis of this approach. In sum, physics-based methods calculate the binding energy by adding up the Van der Waals, electrostatic, hydrogen bonding (which could be included implicitly in electrostatic terms), and desolvation energy terms (Liu and Wang, 2015). The general functional form of physics-based methods is shown in the following equation:
$\Delta G_{\text{binding}}=\Delta E_{\text{vdw}}+\Delta E_{\text{electrostatic}}+(\Delta E_{\text{H-bond}})+\Delta G_{\text{desolvation}}$ (3)
Physics-based methods take advantage of their ab initio nature, where no additional information about the protein and ligand is required other than the binding pose. The interactions can be described using molecular force fields or quantum chemistry calculations.
Physics-based methods for binding affinity prediction consider detailed atomic interactions and capture subtle effects like solvation, entropic contributions, and conformational flexibility. These methods offer mechanistic insights into binding processes, identifying key residues and interactions critical for binding. They can also account for complex cases involving significant ligand and protein conformational changes. When accurately parameterized, these methods can reliably provide energy landscapes for novel compounds, helping to understand the kinetics and thermodynamics of binding (Kuhn et al., 2020; Guterres and Im, 2020; Fu et al., 2022).
However, these methods are computationally intensive and time-consuming, often requiring substantial computational resources and long simulation times (Guterres and Im, 2020). Simulation results are also sensitive to the choice of force fields, simulation parameters, and initial conditions (Fu et al., 2022). Their application is often limited to small to medium-sized systems due to the high computational cost. Additionally, achieving convergence in simulations can be very difficult, which can affect the reliability of binding affinity predictions (Fu et al., 2022).
3.1.2 Empirical methods.
Empirical methods are the most commonly used score-based models, found in packages such as Rosetta and Autodock (Rohl et al., 2004; Trott and Olson, 2010). In addition to the binding energy terms from the previous physics-based methods, they include contributions from empirical chemistry factors, such as hydrophobicity, metal-ligand interactions, entropy effects from steric hindrance, and ligand motifs (Rohl et al., 2004; O’Meara et al., 2015; Park et al., 2016; Alford et al., 2017; Halgren et al., 2004; Friesner et al., 2006; Quiroga and Villarreal, 2016). These methods employ statistical learning algorithms such as multivariate linear regression (MLR) or partial least squares regression (PLS) to parameterize the individual energy terms with weights to estimate the binding affinity. Generally speaking, empirical methods are more flexible and adaptive, allowing users to add customized terms to the scoring function. They adopt a generalized functional form such as:
$Score=E_{bind}+E_{emp}$ (4)
where $E_{bind}$ is binding affinity from the physics model and $E_{emp}$ represents empirical terms that account for interactions and contributions that pure physics fails to capture due to the limitations of force field parameters. The latter additional terms improve the binding affinity prediction based on pure physics models and can be tuned to achieve a better fit for particular interests, such as pharmaceutical importance (Wang et al., 2004b; Friesner et al., 2006). However, it requires empirical knowledge to set up the right scoring functions and training to optimize the weight of each term.
3.1.3 Knowledge-based potential methods.
As the name suggests, knowledge-based potential methods rely on learning from a database of protein-ligand complexes to determine the potential between atom pairs and predict binding affinity. The main assumption behind these methods is that if an atom pair, one from the ligand and one from the protein, appears with a higher frequency than the reference distribution, it might indicate an energetically favorable interaction between the given pair. Thus, the distance-dependent potential of an atom pair can be obtained through inverse Boltzmann analysis based on the measured occurrence frequency of this pair across the entire knowledge base.
In some works, these methods are also referred to as knowledge-based methods (Gohlke et al., 2000) or potential of mean force methods (Su et al., 2009). The general functional form of knowledge-based potential methods is represented by the following equation:
$\text{Score}=\sum_{i\in\text{ligand}}\sum_{j\in\text{protein}}-k_{B}T\ln{\big{(}\frac{\rho_{ij}(r)}{\rho^{*}_{ij}}\big{)}}$ (5)
where $k_{B}$ is the Boltzmann constant, $T$ is the absolute temperature, and $r$ is the distance between atom pairs. The knowledge-based methods are more general, implicitly incorporating effects not fully understood from the structural data. Additionally, they can also incorporate some of the physics-based or empirical terms to enhance performance.
Knowledge-based potential methods have the the following advantages: (1) They are computationally simple. (2) To obtain the distributions of atom-pair occurrences, these methods can take advantage of complex structures without measured binding affinity. Their disadvantages are as follows: (1) The distribution of atom pairs is only relevant to the binding pose rather than binding affinity, which theoretically makes knowledge-based potential methods inferior to other methods. However, in practice, this drawback does not seem to hurt performance, as knowledge-based potential methods are comparable to others. (2) The distribution statistics obtained from the knowledge base are not accurate and should not be assumed to be in a Boltzmann distribution. This means that the potential terms in the functional forms are only approximations.
3.1.4 Hybridization of these three methods
Efforts have been made to bring different categories of scoring functions together to improve performance, blurring the boundaries between these categories. Some physics-based methods use weight parameters derived from regressions to increase performance. Additionally, some knowledge-based potential methods add solvation and entropy terms (Liu and Wang, 2015).
Multiple reviews and comparative studies have been conducted to access the convectional scoring functions. A comparative study of 16 popular conventional scoring functions by Cheng et al. (2009) indicates that no single one consistently outperforms the others in scoring, ranking, and docking. The Pearson correlation between the scores and the experimentally determined binding affinity of the most commonly used scoring functions, such as Glide (Halgren et al., 2004), AutoDock, and Dock, ranges from 0.4 to higher 0.6. This indicates a noticeable correlation and provides some justification for using docking, but it is not sufficient to be applied in virtual screening and drug discovery with high confidence. However, despite the sub-par performance in binding affinity prediction, multiple studies have found that docking software with conventional scoring functions can accurately predict the correct binding pose and conformation (60-80%) (Cheng et al., 2009; Plewczynski et al., 2011; Li et al., 2014). This accuracy is not typically achieved by descriptor-based scoring functions (Xie and Hwang, 2014).
Conventional scoring functions were proposed in the early stages of research on structure-based virtual screening and are widely used in almost all commercial and academic docking software. These scoring functions are still useful and relevant to research on protein-ligand binding today. Most of them are based, to some extent, on the theory of physics and chemistry, which makes them not only reliable and explainable to some degree but also capable of improvement as the theory advances. Cavasotto and Aucar (2020) showed that incorporating a quantum mechanical semi-empirical PM7 Hamiltonian as a scoring function significantly improved the number of compounds correctly screened using molecular docking.
Another advantage of conventional scoring function is that, despite the use of statistical machine learning algorithms, their simple function form and low learning capacity allow them to perform reasonably well without training on too much data. This characteristic was extremely desirable before and during the 2000s, when there were not enough complex structures available for the training of complex learning algorithms and models.
Yet despite these advantages and some recent progress, conventional scoring functions are losing momentum in both research and industry. The disadvantages of conventional scoring functions are mostly rooted in their additive functional forms (Li et al., 2015). First of all, the energy/potential terms in the additive functions are assumed to be independent with each other, which is often not the case, especially for scoring functions based on energy terms (Khamis et al., 2015). Secondly, these functions are not expressive enough for non-linearity, indicating their incapability of complex curve-fitting and conditional branching. Therefore, these scoring functions cannot take advantage of higher-quality data (Ballester et al., 2014) and the growing number of available protein-ligand complex structures. Lastly, even from a theoretical point of view, upon which the conventional methods were usually based, these scoring functions ignored many important aspects of binding, such as the implicit treatment of solvent and protein flexibility (Ballester and Mitchell, 2010a), which became sources of error that are difficult to address. All these drawbacks are related to two factors that are deeply rooted in all conventional scoring functions: (1) the underlying theory of docking is incomplete, and the scoring functions, as a way to approximate the binding affinity, are flawed in many ways (Pantsar and Poso, 2018); (2) the additive functional form neither accurately reflects the physio-chemical process of binding nor is suitable for learning from the increasing number of available protein-ligand complex structures.
3.2 Machine learning-based methods.
The high cost of conventional physics-based approaches, as implemented in tools such as Openeye, leads to growing interest in machine-learning approaches to binding affinity prediction. The scientific community has primarily pursued interaction-free and interaction-based models, as we discuss in the next two subsections. These methodologies are distinguished by whether or not they rely on physical interactions to inform predictions.
3.2.1 interaction-free approach
Interaction-free models infer binding affinity from data without focusing on direct physical interactions. Specifically, these ML-based models consist of two separate parts, each aiming to learn representations from molecule and protein data, including SMILES, protein sequences and graphs. In this approach, the interactions between proteins and ligands are implicitly captured in the latent spaces of their embeddings, which are formed through a neural network that processes their concatenated representations.
Early models in this field used 1D sequence-level representations, such as SMILES for molecules and protein sequences for proteins (Öztürk et al., 2018; Abbasi et al., 2020; Wang et al., 2021; Yang et al., 2021; Zeng et al., 2021; Zhao et al., 2022; Yuan et al., 2022). These early models were innovative but had limitations because they did not include important information about the 3D structures of molecules and proteins. This lack of 3D information often led to less accurate predictions of binding affinities.
With the development of graph neural networks (GNNs), researchers began using more advanced representations. They started by representing molecules as graphs, which better captured the relationships between atoms, while still using linear sequences for proteins (Nguyen et al., 2021; Yang et al., 2022). This change allowed for more accurate modeling of molecular structures and improved the models’ predictions.
As GNNs improved, researchers began representing complex protein 3D structures as graphs too. This approach, where both proteins and molecules are represented as graphs, has greatly enhanced the models’ ability to capture the details of molecular interactions (Somnath et al., 2021; Jiang et al., 2022; Wang et al., 2023; Liu et al., 2024b). By including spatial information, these models give a more complete view of protein-ligand complexes, leading to better predictions of binding affinity.
Additionally, the growing interest in multimodal models in the ML community has led to the development of models that combine different types of data. Researchers have created multimodal models that use both 1D sequences and 3D graphs to represent proteins (Wu et al., 2023; Zhang et al., 2023; Wu et al., 2024). These models take advantage of the strengths of different data types, combining detailed sequence information from 1D sequences with spatial and relational information from 3D graphs. This combination is especially useful for capturing complex interactions that might be missed by using only one type of data.
However, even with these advancements, interaction-free models still face challenges. The inputs to these models often lack detailed interaction information, making it difficult to accurately predict protein-ligand binding affinity. Moreover, GNNs often struggle to capture essential long-range interaction information between proteins and ligands, which is crucial for predicting binding affinity accurately.
Models Protein Repr Molecule Repr Dataset Used
(Öztürk et al., 2018) Sequence Sequence Davis, KIBA
(Abbasi et al., 2020) Sequence Sequence Davis, KIBA, BindingDB
(Wang et al., 2021) Sequence Sequence PDBBind
(Yang et al., 2021) Sequence Sequence Davis, KIBA, Metz
(Zeng et al., 2021) Sequence Sequence Davis, KIBA
(Zhao et al., 2022) Sequence Sequence Davis, KIBA, Metz
(Yuan et al., 2022) Sequence Sequence Davis, KIBA
(Nguyen et al., 2021) Sequence Graph Davis, KIBA
(Yang et al., 2022) Sequence Graph Davis, KIBA, Metz
(Somnath et al., 2021) Graph Graph PDBBind
(Wang et al., 2023) Graph Graph PDBBind
(Jiang et al., 2022) Graph Graph Davis, KIBA
(Wu et al., 2023) Multi (Graph + Seq) Graph PDBBind
(Zhang et al., 2023) Multi (Graph + Seq) Multi (Graph + Seq) Davis, KIBA
(Wu et al., 2024) Multi (Graph + Seq) Graph Davis, KIBA
(Zhang et al., 2023) Multi (Graph + Seq) Multi (Graph + Seq) Davis, KIBA
(Liu et al., 2024b) Multi (Graph + Seq) Graph PDBBind
Table 2: Comparison of Models for Interaction-free Binding Prediction
3.2.2 interaction-based approach
Interaction-based models make predictions based on the 3D structures of complexes and the physical interactions between proteins and ligands. These models employ only the atoms surrounding the interaction/binding pocket to build graphs for prediction. Interaction-based methods are primarily dominated by 3D voxel grids and graphs, relying on 3D CNNs and GNNs, respectively. 3D voxel grid-based methods (Jiménez et al., 2018; Li et al., 2019; Hassan-Harrirou et al., 2020; Jones et al., 2021) use a 3D voxel grid to represent the 3D structures of protein-ligand complexes as input features for CNNs. However, the 3D voxel grid representation has several limitations. First, it has high memory consumption and computational cost due to the sparsity of the voxels, as the protein structure occupies only a small part of the entire grid. Additionally, the sensitivity of 3D grids to rotation negatively impacts prediction results. In contrast, graph-based methods (Jiang et al., 2021; Li et al., 2021; Moon et al., 2022; Yang et al., 2023; Yu et al., 2023) are mostly rotation-invariant, making their graph representations more robust than grid representations. However, interaction-based models are limited to known protein-ligand complex structures, making them less useful compared to interaction-free methods when faced with unknown protein-ligand complexes.
To address these challenges, recent research has focused on developing models that can simultaneously predict high-quality protein-ligand complexes and binding affinity (Lu et al., 2022; Tan et al., 2024). These innovations hold the potential to overcome the limitations of both interaction-free and interaction-based methods by generating accurate predictions even in the absence of known complex structures.
Models Repr Dataset Used
(Jiménez et al., 2018) Voxel PDBbind
(Li et al., 2019) Voxel PDBBind
(Hassan-Harrirou et al., 2020) Voxel PDBBind
(Jones et al., 2021) Voxel PDBBind
(Jiang et al., 2021) Graph PDBBind
(Li et al., 2021) Graph PDBBind, CSAR
(Moon et al., 2022) Graph PDBBind, CASF, CSAR
(Yang et al., 2023) Graph PDBBind, CSAR
(Yu et al., 2023) Graph BindingDB
(Lu et al., 2022) Graph PDBBind
(Tan et al., 2024) Graph PDBBind, CASF
Table 3: Comparison of Models for Interaction-based Binding Prediction
4 Evaluation
Different from most numeric regression problems, the evaluation of binding affinity prediction is much more complex than simple error assessment, which is pointed out repeatedly in multiple comparative studies and benchmarking papers (Cheng et al., 2009; Ashtawy and Mahapatra, 2012; Li et al., 2014; Khamis and Gomaa, 2015; Liu et al., 2017; Su et al., 2018). Evaluation by ranking with decoy ligands and binding poses is necessary to simulate the process of virtual screening, and hereby make sure of the model’s practical usefulness. Cheng et al. (2009) evaluated scoring functions or binding affinity predictors from three perspectives: scoring power (binding affinity prediction), docking power (binding pose prediction), and ranking power (ligands relative ranking prediction). In a more recent study by the same group, Li et al. (2014) enriched the evaluation by adding screening power (discrimination of true binders from decoys) to the set.
In the following subsections, we review the concepts, procedures, and metrics of all four aspects in the evaluation of protein-ligand binding affinity predictors.
4.1 Scoring Power
Scoring power (Su et al., 2020) refers to the ability to produce binding scores that are correlated to the experimentally measured affinities, preferably in a linear fashion. For conventional scoring functions, the scoring power could be measured with Pearson’s correlation coefficient ($R$) or with standard deviation ($SD$):
$R=\frac{\sum(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum{(x_{i}-\bar{x})}^{2}}\sqrt{\sum{(y_{i}-\bar{y})}^{2}}},~{}~{}~{}SD=\sqrt{\frac{\sum{[y_{i}-(ax_{i}+b)]}^{2}}{N-1}}$ (6)
where $x_{i}$ is the prediction for the $i$th complex in the evaluation set, $y_{i}$ is the experimental binding affinity, $a$, $b$ are linear regression factors between scores and binding affinity, and $N$ is the number of samples in the evaluation set.
However, machine learning methods can often estimate the binding affinity directly, instead of producing a score. In these cases, common regression metrics such as MAE, MSE or RMSE can also be used. Still, Pearson’s correlation coefficient ($R$) seems be the most commonly used one for its simplicity and invariance to scaling and unit. There exists other statistical indicators that reflects the linear correlation between the predictions and targets. However, these indicators often correlate with Pearson’s correlation coefficient and do not provide extra information.
Scoring power is the most essential ability of binding affinity models, which means that the evaluation of scoring power is often adequate for studies that focus solely on affinity prediction. However, in most cases, we are also curious about how the models perform in different scenarios or real-life settings, such as docking pose prediction and virtual screening. It is more than common, according to Li et al. (2014), to see a model make excellent predictions for binding affinity, but be incapable of telling the true ligand from decoys, which ultimately renders it useless in drug discovery. Hence other evaluations for docking, ranking, and screening, are required in order to assess the models in a more comprehensive way.
4.2 Docking Power
Docking power refers to the ability that a model can tell the true binding pose from the decoy poses of a given protein-ligand pair. Ideal, the complex with native binding pose should be the one with the highest score or predicted affinity, compared to all the decoys. To implement such evaluation, complexes with decoy binding poses are usually generated with molecular docking program or through molecular dynamics simulation. In the later versions of CASF (Li et al., 2014), they used multiple docking programs such as AutoDock (Huey et al., 2012), LigandFit (Venkatachalam et al., 2003), GOLD (Verdonk et al., 2003), Surflex (Jain, 2003), and FlexX (Schellhammer and Rarey, 2004) for decoy pose generation, to minimize the bias in conformation sampling. The native/true pose is explicitly added into the decoy set so that there is at least one correct binding pose in CASF, according to Cheng et al. (2009). After the scores/affinities for all decoy poses are generated, we compare the pose with the highest score/affinity against the true pose using RSMD, which is calculated with:
$\text{RMSD}=\sqrt{\frac{\sum[{(x_{i}-x^{\prime}_{i})}^{2}+{(y_{i}-y^{\prime}_{i})}^{2}+{(z_{i}-z^{\prime}_{i})}^{2}]}{N}}$ (7)
where $(x_{i},y_{i},z_{i})$ and $(x^{\prime}_{i},y^{\prime}_{i},z^{\prime}_{i})$ are the coordinates of the $i$th atom in the true and predicted decoy poses separately.
Lower RMSD is more desirable because it means that the predicted pose is closer to the native one. Therefore, we consider a model successfully predicted the docking pose for a protein-ligand pair if the RMSD between the true and predicted poses is below a given cutoff. In CASF, the docking power of a model is evaluated by the success rate under different cutoffs ($1.0$, $2.0$, and $3.0$) for RMSD.
One can imagine there are multiple ways to evaluate the docking power. For example, the RMSD between the native and predicted poses can be directly used for assessment. Alternatively, we can simply use the accuracy for predicting the complex in native binding pose with the highest score/affinity. However, due to the resolution and different structure measurement techniques, the native poses obtained experimentally might not be the truly optimized ones. Also, we have to take the flexibility of proteins into account. Therefore, using a static pose directly without a cutoff might cause misleading evaluation results.
Aside from the aforementioned evaluation issue, RMSD, as a metric for distance deviation, has been questioned in multiple independent studies for its drawbacks, such as sensitiveness to the atom ordering, and unawareness of the flexibility of protein structures. Li et al. (2014) used RMSDPM which matches atom pairs between two poses using atom types instead of the atom IDs. Damm and Carlson (2006) proposed wRMSD, a weighted alternative to RMSD that takes into account the flexibility of proteins, which was then recommended by Carlson (2016) as a better alternative to naive RMSD.
4.3 Ranking Power
Ranking power refers to the ability that a model can correctly rank the know ligands by their binding affinity to the given target protein, usually with the provided poses.
In the ideal case, a binding affinity scoring function/predictor should rank these three ligands accordingly, and therefore the evaluation metrics could be one of Spearman’s rank correlation coefficient ($\rho$) or Kendall’s rank correlation coefficient ($\tau$), defined as the following:
$\rho=$ $\frac{\sum(r_{i}-\bar{r})(r^{\prime}_{i}-\bar{r^{\prime}})}{\sqrt{\sum{(r_{i}-\bar{r})}^{2}\sum{(r^{\prime}_{i}-\bar{r^{\prime}})}^{2}}}=1-\frac{6\sum{(r_{i}-r^{\prime}_{i})}^{2}}{n(n^{2}-1)}$ (8)
$\tau=$ $\frac{2}{n(n-1)}\sum_{i<j}\text{sgn}(r_{i}-r_{j})\text{sgn}(r^{\prime}_{i}-r^{\prime}_{j})$ (9)
where $r_{i}$ and $r^{\prime}_{i}$ are the true and predicted rank of the $i$th member in the test set of size $n$.
Besides these two commonly used statistical correlation coefficients, some other metrics are used to assess the rank power. For instance, CASF adapts the predictive index (PI) proposed by Pearlman and Charifson (2001) to measure rank, which places a higher weight on the complex pairs with significant differences in binding affinities. PI is computed as:
$\text{PI}=\frac{\sum_{i<j}[(T_{j}-T_{i})\cdot\text{sgn}(P_{j}-P_{i})]}{\sum_{i<j}|T_{j}-T_{i}|}$ (10)
where $T_{i}$ and $P_{i}$ are the target and predicted binding affinity for the $i$th test sample.
Another example is CI (concordance index) (Gönen and Heller, 2005), used for example by Öztürk et al. (2018) and Öztürk et al. (2019) to measure concordance of target and predicted ordering of binding affinities, computed as:
$\text{CI}=\frac{\sum_{T_{i}<T_{j}}[0.5+0.5\cdot\text{sgn}(P_{j}-P_{i})]}{\sum_{T_{i}<T_{j}}1}$ (11)
The latest version of CASF provides a test set with 57 target proteins, each with five ligands that binds the target with distinctively different affinities, resulting in 285 protein-ligand complexes. The standard metrics include Spearman’s rank correlation coefficient ($\rho$), Kendall’s rank correlation coefficient ($\tau$), or Predictive Index (PI).
4.4 Screening Power
Screening power (Guedes et al., 2018) refers to the ability of a model to tell a true ligand binder from decoy molecules for a given protein target. For the latest version of CASF, there are 57 target proteins and 285 ligands. Each protein has five unique positive ligands, and the other 270 ligands are used as decoy molecules. For each protein and each decoy molecule, the authors have collected 100 potential binding poses from various softwares. The tested scoring function or binding affinity predictor will be applied to every pose, and the highest score/affinity will be considered as the score/affinity of the this protein-ligand pair. And finally, the screening power of the model is tested by ranking all the ligands in the descending order of their score/affinity and see if all five true ligands are ranked in the top, which is characterized with enhancement factor (EF), computed as:
$\text{EF}_{\alpha}=\frac{\textit{NTB}_{\alpha}}{\alpha\cdot\textit{NTB}_{\text{total}}}$ (12)
where $\alpha$ is an arbitrary percentage threshold, typically 1%, 5%, or 10%; $\textit{NTB}_{\alpha}$ is the number of true binders found among the top $\alpha$ candidates; and $\textit{NTB}_{\text{total}}$ is the total number of true binders for a single target protein, which in CASF’s case is 5. A major problem with CASF’s approach is that they assumed that random ligands in the dataset do not bind with a target protein, which is often not the case. In fact, Su et al. (2018) report that 21 of the 57 target proteins have more than five binding ligands in the dataset according to known binding data in ChEMBL, which makes CASF’s evaluation for screening power rather questionable. Moreover, using random molecules as non-binders in a decoy set is a common practice in the research of protein-ligand binding affinity and drug-target interaction, which could potentially hinder the learning and validation for the models.
Among all four evaluation aspects, scoring power, on the surface level, seems to be the dominant factor over the other three. Essentially, docking power, ranking power, and screening power are evaluated with the scores/affinities predicted from the model. Thus, it is completely reasonable to assume that a binding affinity predictor with good scoring power might also perform well on the other three evaluations. However, it seems that in the literature such as (Cheng et al., 2009), it is simply not the case. For conventional methods, we usually found that ranking is a much harder task and for machine learning based predictors, despite their impressive results for scoring power, they seem to perform poorly on docking, which renders the whole category of methods unconvincing.
5 Conclusion
We have provided an extensive review of binding affinity prediction methodologies, tracing their evolution from conventional approaches to the latest machine learning (ML) and deep learning (DL) techniques. Binding affinity, which refers to the strength of interaction between a small molecule (like a drug) and a target protein, is crucial in drug discovery and other areas of bioinformatics. Our review has highlighted the increasing application of ML and DL methods in this domain, driven by the growing availability of protein and small molecule data.
We categorized binding affinity prediction methods into three main types: conventional methods, traditional machine-learning-based methods, and deep-learning-based methods. Conventional methods rely heavily on established chemical knowledge and physics-based calculations, often resulting in rigid models that perform well only in specific cases. In contrast, traditional ML methods improve prediction accuracy by leveraging human-engineered features, while ML methods, which have gained prominence more recently, aim to automate feature extraction and can potentially outperform other methods as more data becomes available.
We also discussed challenges in binding affinity prediction, such as the limitations of available datasets, which often contain biased or insufficient data, and the inherent complexity of protein-ligand interactions that make accurate predictions difficult. We emphasize the need for more robust datasets and models that can handle the dynamic nature of protein-ligand interactions.
Our review also leads us to identify five important areas for further research, as follows. 1) Overcome limitations in existing datasets. Current datasets are often biased towards protein-ligand pairs with favorable binding constants and correct poses, while complexes with low affinity or failed bindings are underrepresented. Future research could focus on creating more balanced datasets that better represent the full spectrum of binding affinities. 2) Integrate physics-based models with ML/DL. There is potential for hybrid models that combine the theoretical rigor of physics-based methods with the predictive power of machine learning. Research in this area could explore how best to integrate these approaches to improve prediction accuracy. 3) Better handle protein flexibility. Most existing models treat proteins as rigid entities, which is a significant simplification. Developing methods that account for protein flexibility, as well as different conformation of ligand, during binding could lead to more accurate predictions. 4) Improve evaluation metrics. Current evaluation metrics, such as Pearson correlation for scoring power and RMSD for docking power, have limitations. Research could explore alternative metrics that better capture the nuances of binding affinity predictions, particularly in real-world applications like drug discovery. 5) Multimodal data integration. With the growing interest in multimodal models, future research could focus on integrating diverse data types (e.g., sequence data, 3D structures, interaction networks) to capture complex interactions more effectively. Research in each of these areas promises to advance the field of binding affinity prediction, potentially leading to more accurate and reliable tools for drug discovery and other applications.